OpenStudy (sissyedgar):

Solve for m. 4+|7−m|=5 A. m = 6 or m = 8 B. m = 6 or m=−8m=−8 C. m = 16 or m=−2m=−2 D. m = 2 or m = 16

OpenStudy (3mar):

How can you expand the absolute value of |7-m|? Do you know how?

OpenStudy (otherworldly):

-4 on both sides of the equal sign then do absolute value

OpenStudy (sissyedgar):

Well.. I was thinking... The answer is probably A. because when I do the math I get m = 8 A. is the only one with 8 as an option.

alones (alones):

A would be right. since than you get \(\ 6\)

OpenStudy (sissyedgar):

Yay. Thx!

alones (alones):

\(\ m+7=1\) Subtract \(\ 7\) from \(\ −m+7−7=1−7\) \(\ −m=−6\) Divide both sides by \(\ -1\) \(\ −m−1=−6−1\) So we'll get \(\ m=6\)

OpenStudy (rcp031):

4+|7−m|=5 |7−m|=1 7-m=-1 & 7-m=1 m=8 or m=6

OpenStudy (rcp031):

Consider |x|<2. Absolute value is defined as distance from zero. Another way to read this inequality would be the distance from zero is less than 2. So on a number line we will shade all points that are less than 2 units away from zero.

OpenStudy (sissyedgar):

Thankz guyz!